14.64 Paper: Working

Jim Grayson (and others from The Making)

Chris Laas

May 1, 2000

Jim Grayson is a spot welder at the Ford assembly plant on the South Side of Chicago. He stands at an assembly line post, from 3:30 PM to midnight every weeknight, and welds pieces of metal together by pressing buttons on a bulky and heavy welding machine: one button for high voltage, and one button for low voltage. Staying always within a two or three-foot area, he works on about four hundred cars a day.

The company is always trying to improve efficiency; manufacturing process engineers make detailed measurements of the time it takes to perform any task, and especially of any time wasted by, for example, walking away from an assembly line post to get more stock. If any inefficiencies in the line are detected, they are corrected immediately: in this example, stock is brought directly to the post in a truck. In the General Motors plant in

Chris Laas  14.64  5/1/2000  2

Lordsdown, Ohio, twenty-two welding robots called “unimates” pushed production from sixty cars an hour to one hundred cars an hour, while replacing about two hundred workers. The firms substitute capital, in the form of information and machinery, for labor, as long as it is more productive for the price; the process is slower than simply hiring more workers, but the result is more production efficiency, and ultimately more profit for the firm.

Jim Grayson doesn’t enjoy his job. The work is uncomfortable: he stands up all day, pushes around heavy machinery in a repetitive way, and can only go to the bathroom during the two short breaks in his shift. “Plant pollution,” such as carbon monoxide, disturbs him, and the plant is full of immediate hazards such as sparks and smoke and heavy machinery with moving parts. The head foreman at Grayson’s plant knows that cuts and other injuries are everyday occurrences; Grayson feels that the night shift medic department is inadequate. Most importantly, Grayson has little variety in his work, cannot talk to the other workers while at his post, and, altogether, is bored.

So, why does Grayson stay in this job? There’s a lot of evidence that he is getting a compensating differential in his wage, relative to what he would be getting elsewhere. Phil Stallings, in the same job as Jim, gets $4.32 an hour, and thinks “that’s real good money now;” the plant manager, Tom

Chris Laas  14.64  5/1/2000  3

Brand, notes that Ford pays “top dollar” and that the workers get twelve paid vacation days a year. Brand also expounds with pride upon his concern for the safety of his workers: he describes measures including a doctor, four nurses, prescription safety glasses, non-slip shoes, hearing protection, and powerful cooling fans. There’s no doubt that, without these measures, the compensating wage differential would have to be significantly higher in order to hire sufficient labor.

The best evidence that Grayson receives a compensating wage differential in this job is that he declares that he would be willing to accept a reduced wage if he could get a job in “industry”, by which he means a desk job in which he could talk with other workers and do nonrepetitive work. To this end, he is enrolled in a part-time Business Administration degree program at Roosevelt University; he wants a job with “status,” which would clearly give him some utility.

Grayson’s decision to invest in education could be explained under either the human capital or the signaling model, since the jobs he wants require some general and specialized training but probably not a college education. There is some evidence that it fits more with the signaling model, however: Grayson applied to a junior accountant position for which he was qualified, but he feels he was turned down because he was Black. The fact that he

Chris Laas  14.64  5/1/2000  4

continues to invest significant money, time, and energy into college implies that he believes that better credentials (signals) will improve his odds in the job market, despite the fact that he feels that he already has enough human capital to enter the market for “desk jobs with status.”

It is also notable that, while Grayson is investing heavily in education with the expectation of greater future utility, Phil Stallings (who is very similar to Grayson in most other ways) is not investing in formal eduction at all; his ambition is to become a utility worker for the plant. To some small degree, Stallings is investing in specialized training, and he is signaling his productivity via hard work, to encourage the company to promote him. It’s quite clear that Grayson is investing more heavily in his future than Stallings, however; in economic terms, this is no doubt due to Grayson having a lower discount rate than Stallings. Grayson is more future-oriented, and Stallings is more present-oriented.

Young workers are more mobile, and Gary Bryner, the union president at the General Motors plant in Lordstown, comments that they are more likely to be taking classes part-time; this is certainly consistent with our model of education, since older workers have a shorter payoff period, limited, in this case, by Ford’s thirty-and-out policy. In general, the old have more stable lifestyles and are less likely to rock the boat, since they look

Chris Laas  14.64  5/1/2000  5

forward to their thirty-and-out retirement packages coming up; the old also have a more difficult time finding a new job. Thus, they are less likely to leave the company, and in this way, the firm retains the workers with more experience (specialized on-the-job training) and puts them in higher positions, such as utility worker or foreman. Nevertheless, old workers without specialized training are less productive than young workers, and so the company statistically discriminates against them; the firm hires more young people with low wages and high productivity.

There’s evidence that Blacks are discriminated against at Ford: for example, Grayson feels that the foremen, mostly Southern Whites, unfairly single out Blacks. Bryner points out that older workers call each other “nigger” and “honkie”. The plant manager admits that although the plant workers are forty-five percent Black, only twenty-five percent of salaried workers are Black. There doesn’t seem to be any evidence for statistical discrimination here, though, because there’s no sign that Blacks are less productive that Whites in any part of the plant. However, taste-based discrimination on the part of employees provides a good explanation for these effects: Whites in the company management have no problem with Blacks working in jobs of lower status, but these Whites have lower personal utilities if Blacks share the same rank as them. The predominantly white foremen

Chris Laas  14.64  5/1/2000  6

provide strong support for this hypothesis. However, this model does not explain the relatively high level of integration on the lowest level, the assembly line; this could be explained by social reasons (the young unskilled workers don’t mind integration as much as the older leadership or the more Southern foremen) or could be explained by compensating differentials if it were found that White workers in a completely segregated plant earned less than workers in this Ford plant.

Another interesting point is that, until recently, women did not work on the assembly line at all; the work was too physically demanding. In a way, this was an extreme case of a kind of statistical discrimination, in that women, in general, had very low productivity on the assembly line, and thus were never hired for that kind of work. Of course, probably very few women ever applied for the job; and, regardless of gender, an employer can easily estimate the size of the muscles on a prospective employee. Thus, the label “statistical discrimination” is somewhat questionable for this effect, since it has less to do with statistically educated guesses and more to do with an appraisal of actual applicants.

One thread that runs through all the essays is the issue of how the firm prevents shirking. Grayson points out that absenteeism is rampant: workers are experts at producing fake doctor’s notes and the like to have

Chris Laas  14.64  5/1/2000  7

extra leisure time on Mondays. When workers like Grayson do not take pleasure in their work, the company must find ways to keep assembly line workers producing. One way is the use of efficiency wages; this, along with compensating differentials, helps to explain why the wages at the plant were relatively high. The monitoring strategy of the plant, hiring thirteen foremen to supervise the workers, is expensive, but not so costly as lost productivity due to shirking workers; the close monitoring allows the firm to decrease the no-shirk premium.

The primary punishment for shirking is to be sent to the “Green House”, where one would be either reprimanded or fired; the promise of promotion (status) and pay increases (the utility man makes $4.49 \frac{1}{2} per hour, compared to Grayson’s $4.32 per hour), along with the touted thirty-and-out retirement plan, provide incentives to avoid this fate. Interestingly, however, the company does not usually fire workers permanently, but instead docks them a few days’ pay and hires them again. This is the best strategy for the firm, because it would no doubt have a strong negative effect on a worker’s chances of promotion, so would still remain a disincentive for shirking; but would alleviate some of the hiring and training costs of the firm, since it could re-hire workers who already knew the trade. This theory is supported by the high turnover rate at the plant — at one point, the firm hired 125

Chris Laas  14.64  5/1/2000  8

people per week, which is a 2.7% per week turnover rate. However, Brand, the plant manager, notes that during a recession, the turnover rate (and shirking) becomes much less of a problem, due to the low employment rate; this, again, lends credence to the theory.

Jim Grayson complains that he is treated like a robot; in economic terms, however, he is almost as predictable as one. The manufacturing-plant situation in which he finds himself is very well studied, and the models of productivity, discrimination, wage determination, and education explain the qualitative aspects quite well. It is likely, however, that a quantitative analysis would reveal some portions of the effects, especially discrimination and education, which are not as easily explained by the simple models; these would be attributed to social effects which were not incorporated into the models. It is also clear that, in some aspects, the factory in which Grayson works is not in efficient equilibrium; for example, some of the working conditions which Grayson (and the union) would prefer might, in fact, be possible for the factory to create without significantly decreasing productivity, but the plant manager seems oblivious to such concerns. A more detailed model, taking into account the effects of incomplete information and unbalanced bargaining, might be able to address more of these issues, and create a more complete picture of the microeconomy of this firm.